OpenStudy (anonymous):

I NEED HELP PLEASE...I'LL AWARD MEDAL Find the measure of the central angle when the sector area is 169.56 and the radius is 9.

OpenStudy (anonymous):

@Melodysim

OpenStudy (anonymous):

Can Somebody Please Help Me???

OpenStudy (here_to_help15):

here i will give a site ok but i will get a medal correct?

OpenStudy (anonymous):

Area of a circular sector goes like this A=(1/2)(r^2)(θ) where A is the area of a sector, r is the radius, and θ is the Central angle solving for θ: multiply the equation to 2 so that we can cancel 1/2 2A=(r^2)(θ) divide both sides by r^2 θ=(2A)/(r^2)

OpenStudy (anonymous):

wow lol @Here_to_Help15 you are really obsessed with medals ^^

OpenStudy (anonymous):

so plug in for each value. would you want to do that? @gswag98

OpenStudy (anonymous):

i got the answer it was 240

OpenStudy (anonymous):

yuppp that is it!!! YOU DID IT!! YOU ARE AWESOME!!

OpenStudy (anonymous):

Thanks for medal!!

OpenStudy (anonymous):

You're Welcome and Thank You For The Help @Melodysim

OpenStudy (anonymous):

thanks